OpenStudy (anonymous):

Please help find the x, y, and z intercepts. 2x-6y-3z=12

OpenStudy (tkhunny):

Divide by 12 \(\dfrac{x}{6}+\dfrac{y}{-2}+\dfrac{z}{-4} = 1\) Oddly, few are aware of the Intercept Form. It is a beautiful thing.

OpenStudy (anonymous):

Ohhh, I was missing the negative signs. Thank you!

OpenStudy (tkhunny):

Here's a plan... Don't miss those negative signs!! :-)